# Datasets, Documents, and Repetitions: The Practicalities of Unequal Data Quality

### v1

@@ 1 Introduction
Scaling data and compute has been the winning recipe for producing state of the art machine learning models. As costs increase with diminishing returns, recent work often point to high quality data as a key factor for high model performance (Team et al., 2023; Dubey et al., 2024). One method of obtaining high quality data is through data filtering. However, as frontier level large language models (LLMs) scale in compute through increases in both parameter count and token count, there are concerns that we will run out of training data. Data filtering reduces the dataset size, creating a tension between data quality and data quantity. It is currently unclear whether the performance gains from data filtering can be sustained at scale through repetitions as an attempt to compensate for the reduction in dataset size.
@@
Recent work has studied data filtering as a means of efficiently training relatively performant models with limited compute (Gadre et al., 2023; Li et al., 2024). However, it is unclear if the findings are practical for training larger models. Unfortunately, existing frontier level LLMs lack transparency when it comes to their data filtering pipelines (Team et al., 2023; Dubey et al., 2024; Achiam et al., 2023; Liu et al., 2024). This gap in knowledge adds to the concerns about the practicality of data filtering at scale.
@@
Understanding the interaction between data filtering and scaling would be made easier if we better understood what it means to be a better dataset. Currently the quality of a dataset is determined by the performance of a model trained on it as evaluated by standardized benchmarks. However, there are other traits we desire in data beyond raw downstream model performance, such as repeatability and performance over replacement. Developing a better understanding of these traits would allow for more principled construction of better datasets.
@@
Figure 1: Repeating the aggressively filtered dataset DCLM-baseline for up to ten epochs consistently outperforms training on a single epoch of the ten times larger RefinedWeb supserset (cyan)—provided that we adapt the weight decay for the high-repetition runs. On the left, we show that this result appears consistent across compute budgets. On the right we show that, at the largest compute budget tested, adapting the weight decay as a function of repetition allows for a significantly better result versus training on the superset. Results are evaluated on the centered core metric from DCLM, which is a normalized average over 22 tasks. We also include a MMLU version of the right side in Appendix C.
@@
In what follows we will compare DCLM-baseline filtering (Li et al., 2024) with a superset comprising the strictly looser RefinedWeb filtering (Penedo et al., 2023), as well as the commonly studied C4 filtering (Raffel et al., 2020). We find that while better datasets are not necessarily more repeatable, the interaction between dataset performance and repeatability depends on the compute allocation used to train the model. Furthermore, we study practical considerations for scaling up aggressively filtered datasets by repeating them for multiple epochs to account for the reduced amount of tokens available. Datasets can be made more repeatable through regularizing factors, such as increasing weight decay and increasing tokens per parameter. We find that when training on the same number of total tokens, using ten epochs of the aggressively filtered DCLM-baseline can outperform a single epoch of its ten times larger superset—RefinedWeb.
@@
Next, we investigate how to create better datasets through analyzing document duplicates within the dataset. We find that web data filtering results in a bias towards documents that are more often partially duplicated, which was also reported in (Penedo et al., 2024). Deduplicating the resulting dataset will therefore reduce the precision of the overall filtering step. We suggest simple methods that use a quality classifier to manipulate the individual counts of documents to improve dataset quality given a desired token budget, and show that this outperforms using only duplicate counts as in standard deduplication methods.
@@
Figure 2: C4 perplexity as a function of training dataset and repetition count. Over-repeating data leads to diminishing returns irrespective of the training dataset chosen. In the left plot we show that heavy repetition results in a similar degradation in validation loss for all datasets tested. The right hand side shows that this effect holds at larger compute budgets for similarly over-trained models on the DCLM dataset.
@@ 2 Related work
We investigate whether data filtering is justifiable when it requires repeating to make up for the decrease in tokens. Recent works such as DCLM (Li et al., 2024) and FineWeb-edu (Penedo et al., 2024) have shown how data filtering can greatly improve performance. However, these datasets only retain a few trillion tokens, which may not be enough to train state of the art models unless trained for multiple epochs. Our study thus takes into account this scenario where aggressively filtered datasets must be repeated, and compares them with larger but lower quality datasets.
@@
We study the interaction between repeating datasets and datasets of different quality. We build off Muennighoff et al. (2023), which studies training models that are data constrained. They find that for a fixed compute budget, data can be repeated four times before encountering significant diminishing returns compared to training on unique tokens. We differ from this work by looking at behavior across datasets of different quality, varying tokens per parameter and regularization settings, and using downstream evaluation accuracy instead of test loss.
Goyal et al. (2024) studies data filtering for CLIP models and argues that data filtering methods must take into account the computational budget. They find that training a ViT-B/32 for 640M samples on 12.8B LAION filtered samples does worse than 128B unfiltered samples, and that CLIP models trained on large amounts of compute should use less aggressive filtering.
@@
Explicitly repeating tokens can lead to different optimal compute allocations. Hoffmann et al. (2022) studies optimal allocation of parameters and tokens for a giving training computational budget. Their suggested ratio for parameters to tokens is considered Chinchilla optimal, while training for more or less tokens can be considered overtraining or undertraining relative to that model size. Muennighoff et al. (2023) find that under the assumption of repeated data, overtraining is more efficient than staying at Chinchilla optimal.
@@
Prior work has shown that training on deduplicated data is desirable because it improves performance (Lee et al., 2022), reduces privacy risks (Kandpal et al., 2022), and reduces memorization (Carlini et al., 2023). However, more recent works like DCLM and FineWeb-edu use sharded deduplication for both engineering and performance reasons. TxT360 (Liu et al., 2025) takes a slightly different approach, using global deduplication followed by upsampling by the natural distribution to improve performance. We build on existing deduplication methods, and move toward upsampling based on quality metrics on a document level.
@@ 3 Repeating Filtered Datasets
In parallel to LLMs training on larger datasets, researchers are also focusing on filtering data to improve quality. Together, these factors accelerate concerns about running out of training data. In the earlier days of training LLMs, common practice was to train LLMs on a single epoch of data. This is in stark contrast to state of the art vision models, where practitioners would repeat data for many epochs. Though single epoch training may have benefits such as preventing memorization (Lee et al., 2022; Carlini et al., 2023), in this work we study data repetition across datasets from a purely performance perspective. Muennighoff et al. (2023) studied the repeatability of datasets in a data-constrained setting and found that datasets can be repeated for four epochs before noticing diminishing returns compared to training on fresh data. Building off this, we examine datasets of varying quality created through data filtering to understand whether these data filtering approaches affect repeatability. In this section we study datasets “as is” and treat them as a single unit rather than a collection of documents. We use original C4, DCLM’s reproduction of RefinedWeb (we will refer to this as RefinedWeb and RW v2 interchangeably), and DCLM-baseline. See Appendix I to compare datasets “as is” with datasets that are deduplicated.
@@
In Figure 2 we study a scenario similar to the findings in Muennighoff et al. (2023), but instead start with Chinchilla optimal unique tokens for 1B parameter models. On the left side, we observe that C4 validation perplexity eventually sees diminishing returns when training on C4 after repeating. In addition, we also train on higher quality datasets DCLM-baseline and RefinedWeb, and find that they have similar repeatability behavior when evaluating on C4 validation perplexity. In this setting we start with a Chinchilla optimal compute allocation and then repeat epochs by fixing the model size while increasing the token count, and compare this to datasets with more available tokens. On the right side we vary model size, but maintain the same dataset sizes that we repeat. This figure suggests that increasing model size while maintaining dataset size decreases repeatability, but there is not enough information to make a conclusion so we will explore this further later in the section. When we move to evaluating downstream accuracy across a 22 task subset from LLM Foundry (referred to as core centered metric in DCLM), we see in Figure 3 that quantifying repeatability is even less clear. It appears that when comparing using downstream accuracy, repeatability is more difficult to measure, and it is unclear if repeating meets the same diminishing returns seen when evaluating on C4 perplexity. What is common between the two evaluation settings is that datasets that achieve better performance do not seem more repeatable than inferior datasets when examining diminishing returns.
Although better performing datasets are not more repeatable by preventing diminishing returns, in the overtraining setting (Figure 3) better datasets repeated multiple times clearly maintain better performance than an inferior dataset that is not repeated. However, it is important to note that overtraining is practical only in certain settings with priorities such as minimizing inference cost, and there is an assumption in Figure 3 that there is at least enough tokens to train for Chinchilla optimal. This assumption does not always hold when LLMs are scaled up, especially when data filtering is involved.
@@
In contrast to repeating datasets by training for longer, we next examine fixing the token budget and repeating by decreasing the number of unique tokens. This scenario forces us to examine cases where there are fewer unique tokens available than Chinchilla optimal. In Figure 4 we compare for a fixed model size and total token budget how performance varies as you reduce the number of unique tokens available, which causes an increase in the number of repeats for which the model sees the dataset. When there is no longer at least the Chinchilla optimal number of tokens available, repeating the DCLM-baseline dataset with too few unique tokens can lead to worse performance than training for a single epoch on its superset of RefinedWeb filtered data. On the other hand, compute allocations with high tokens per parameter seem to benefit more from higher quality datasets, since DCLM can be repeated more times to equalize performance with unrepeated RefinedWeb. Tokens per parameter is the main factor, as increasing model size while maintaining token count decreases DCLM repeatability to equalize with RefinedWeb, and increasing token count while maintaining tokens per parameter does not consistently increase DCLM repeatability to equalize performance with unrepeated RefinedWeb (Figure1, left). This would also explain why unrepeated RefinedWeb does worse than heavily repeated DCLM in Figure3. But there are increasing diminishing returns to performance with respect to compute when increasing tokens per parameter, so there still needs to be a balance when doing data filtering.
@@
We can mitigate the performance degradation caused by repeating data with regularization in the form of increasing weight decay. In Table 1, we see that the optimal weight decay value increases as we increase the number of repeats. A reasonable schedule is to scale weight decay by roughly the square root of the number of repeats. With the optimal weight decay values, we see in Figure 1 that training for ten epochs of DCLM-baseline outperforms training on RefinedWeb filtered data for a single epoch. This is notable because DCLM-baseline takes the top 10% of documents of RefinedWeb as scored by a FastText classifier. However, it is important to note that this does not mean DCLM-baseline is a strictly better dataset than RefinedWeb filtering. When data is a much more constrained resource than compute, it is likely that diminishing returns from heavily repeating DCLM-baseline will perform worse than repeating RefinedWeb filtered data a couple of times.
@@ 4 Repeating Documents To Create Better Datasets
Duplicate aware subsample 34.1 25.1
@@
We have established that, even as compute budgets increase, models can remain dataset repetition resilient given the right training recipe. We now investigate whether high quality data repetition may be best implemented at a document level, as opposed to increased epochs on the entire dataset. To better understand this question, we dive into DCLM-baseline, which contains many duplicates (fuzzy and exact) because it uses sharded deduplication due to engineering constraints. Penedo et al. (2024) also uses sharded deduplication, observing that global deduplication does not always improve performance. This is in contrast to prior standard practices of deduplication, which recommended global deduplication and training for a single epoch on the resulting dataset.
The performance gains from data filtering clearly demonstrates that not all documents are of equal quality. We can also observe this by comparing subsampling strategies. The standard strategy is uniform subsampling, which results in a dataset that is diverse with very few duplicates, while more likely keeping documents that have more duplicates. A second strategy is to run minhash-based global fuzzy deduplication before subsampling down to the desired token count. Compared to uniform subsampling, incorporating global fuzzy deduplication treats documents with different duplicate counts equally, but both strategies should have few or no duplicates in the resulting subset if subsampling significantly. Lastly, we can mimic the duplication profile of the original dataset by using duplicate aware subsampling. For each group of duplicates, this strategy keeps or removes the entire group with probability equivalent to the duplicate count normalized by the desired dataset reduction. This strategy results in a subset with a significant number of duplicates compared to the other subsampling strategies.
@@
Figure 5: Examples of strategies for count manipulation.
In Table 2 we observe that duplicate aware subsampling performs the worst, suggesting that keeping a significant number of duplicates hurts performance. Uniform subsampling does significantly better than global deduplication before subsampling, which shows that documents in DCLM-baseline with high duplicate count are likely to be of higher quality. This is further supported by the fact that keeping unique copies for documents with greater than seven duplicates achieves competitive performance to uniform; however, this removes over 96% of the documents in DCLM-baseline.
@@
Table 3: 7B models trained on DCLM-baseline for 138B tokens using different document count manipulation approaches. Figure 5 visualizes these strategies. Ensemble ranks documents by the worse rank between FastText score and duplication count.
@@
Table 4: 7B models trained for 138B tokens by doing count manipulation on a duplicate aware subsampled 280B tokens from DCLM-baseline. In this setting, the greedy yet diverse strategy sees each document approximately 3 times. Metric used is FastText score. Note that using duplication count as the metric would be ineffective here because there are not enough unique documents.
@@
The results in Table 3 are unusual in that we are taking an already highly filtered dataset and subsampling even further. Naturally, there are enough high quality tokens to train without repeats, hence the success of the greedy yet diverse strategy. However, this becomes impractical as we scale and there are no longer enough unique tokens to support this strategy. Therefore we now investigate the setting where the dataset we are trying to manipulate is closer in size to the desired token count. We duplicate aware subsample DCLM-baseline down to 280B tokens and apply count manipulation to produce 138B tokens for training a 7B model. As we scale back up, this would allow us to use the strategy on the entire DCLM-baseline to create a dataset close to two trillion tokens. In Table 4, we see that the best strategy is applying a linear function to balance the tradeoff between exploring diverse documents and exploiting high quality documents. Contrasting the best strategy in the two settings suggests that using a good count manipulation strategy is more important in the setting where data is constrained.
@@
Combining duplicate aware subsampling with count manipulation is a powerful tool for studying datasets. Increasing dataset size typically refers to obtaining new tokens of the same quality. DCLM does this by increasing the number of Common Crawl WARC files at the same rate as increasing the training token budget. However, as seen in Figure 6, when the number of WARC files increases, so does the number of duplicates. We find that DCLM-baseline constructed from the DCLM 7B_2x pool contains 33% fuzzy duplicates, but DCLM-baseline constructed from 20 times more WARC files contains 83% fuzzy duplicates. By uniformly subsampling to run smaller scale experiments, this experiment design uses both new data and repetitions when scaling back up. Additionally, it is impractical to deduplicate unfiltered data due to engineering constraints caused by the amount of data, as well as design decisions like what type of deduplication to use. Duplicate aware subsampling gives us the tools to design experiments where data scaling happens through providing new data of equal quality.

### v2

@@ 1 Introduction
Scaling data and compute has been the winning recipe for producing state of the art machine learning models. As costs increase with diminishing returns, recent work often points to high-quality data as a key factor for high model performance [30, 6]. One method of obtaining high quality data is through data filtering. However, as frontier level large language models (LLMs) scale in compute through increases in both parameter count and token count, there are concerns that we will run out of training data. Data filtering reduces the dataset size, creating a tension between data quality and data quantity. It is currently unclear whether the performance gains from data filtering can be sustained at scale through repetitions as an attempt to compensate for the reduction in dataset size.
@@
Recent work has studied data filtering as a means of efficiently training strong models with limited compute [8, 15]. However, it is unclear if the findings are practical for training larger models. Unfortunately, existing frontier level LLMs lack transparency when it comes to their data filtering pipelines [30, 6, 1, 16]. This gap in knowledge adds to concerns about the practicality of data filtering on a scale.
@@
Understanding the interaction between data filtering and scaling would be made easier if we better understood what it means to be a better dataset. Currently, the quality of a dataset is determined by using standardized benchmarks to evaluate the performance of a model trained on it. However, there are other traits we desire in the data beyond the performance of the raw downstream model, such as repeatability and performance over replacement. Developing a better understanding of these traits would allow for a more principled construction of better datasets.
@@
Figure 1: Repeating the aggressively filtered dataset DCLM-baseline for up to ten epochs consistently outperforms training on a single epoch of the ten times larger RefinedWeb supserset (cyan)—provided that we adapt the weight decay for the high-repetition runs. On the left, we show that this result appears consistent across compute budgets (model sizes of 1B, 3B, 7B, and 12.6B). On the right we show that, at the largest compute budget tested (12.6B parameters, 252B total tokens seen), adapting the weight decay as a function of repetition allows for a significantly better result versus training on the superset. Results are evaluated on the centered core metric from DCLM, which is a normalized average over 22 tasks. We also include a MMLU version of the right side in Appendix C.
@@
In what follows, in Section 3 we study how datasets of different quality behave under repetition and explore how to improve multi-epoch performance. We compare DCLM-baseline filtering [15] with a superset comprising the strictly looser RefinedWeb filtering [21], as well as the commonly studied C4 dataset [23]. We find that while better datasets are not necessarily more repeatable, the interaction between dataset performance and repeatability depends on the compute allocation used to train the model. Furthermore, we study practical considerations for scaling up aggressively filtered datasets by repeating them for multiple epochs to account for the reduced number of tokens available. Datasets can be made more repeatable through regularizing factors, such as increasing weight decay and increasing tokens per parameter. We find that when training on the same number of total tokens, using ten epochs of the aggressively filtered DCLM-baseline can outperform a single epoch of its ten times larger superset—RefinedWeb.
@@
Next, in Section 4 we investigate how to create better datasets by analyzing duplicate documents within the dataset. We find that web data filtering results in a bias towards documents that are more often partially duplicated, which was also reported in Penedo et al. [22]. Deduplicating the resulting dataset will therefore reduce the precision of the overall filtering step. We suggest simple methods that use a quality classifier to manipulate the individual counts of documents to improve dataset quality given a desired token budget, and show that this outperforms using only duplicate counts as in standard deduplication methods.
@@
Figure 2: C4 perplexity as a function of training dataset and repetition count. Over-repeating data leads to diminishing returns irrespective of the training dataset chosen. For each dataset, we vary the number of unique tokens available, and then vary the training token budget. In the left plot we train 1B parameter models to show that heavy repetition results in a similar degradation in validation loss for all datasets tested. The right hand side shows that this effect holds at larger compute budgets for similarly over-trained models on the DCLM dataset.
@@ 2 Related work
We investigate whether data filtering is justifiable when it requires repeating to compensate for the decrease in tokens. Recent works such as DataComp-LM (DCLM) [15] and FineWeb-edu [22] have shown how data filtering can greatly improve performance. However, these datasets only retain a few trillion tokens, which may not be enough to train state-of-the-art models unless trained for multiple epochs. Our study thus takes into account this scenario, where aggressively filtered datasets must be repeated, and compares them with larger but lower quality datasets.
@@
We build on the work in Muennighoff et al. [19], which studies training models that are data-constrained. They find that for a fixed compute budget, web-crawl data can be repeated four times before encountering significant diminishing returns compared to training on unique tokens. We will refer to this as repeatability, which is measured for a fixed training token budget by comparing the performance of repeating a subset for multiple epochs with that of a single epoch. Unlike this work, we examine the repeatability of crawl-data from two new angles: 1. whether dataset quality affects repetition resilience; 2. if regularization techniques can dampen the impact of over-repetition.
Goyal et al. [9] studies data filtering for CLIP models and argues that data filtering methods must take into account the computational budget. They find that training a ViT-B/32 for 640M samples on 12.8M LAION filtered samples is worse than 128M unfiltered samples, and that CLIP models trained on large amounts of compute should use less aggressive filtering. Although this differs from our findings, there are significant experimental differences due to modality and filtering techniques.
@@
Explicitly repeating tokens can lead to different optimal compute allocations. Hoffmann et al. [11] studies optimal allocation of parameters and tokens for a training computational budget. Their suggested ratio for parameters to tokens is called Chinchilla optimal, while training for more or less tokens can be considered overtraining or undertraining relative to that model size. Muennighoff et al. [19] find that for repeated data, overtraining is more efficient than staying at Chinchilla optimal.
@@
Prior work has shown that training on deduplicated data is desirable because it improves performance [13], reduces privacy risks [12], and reduces memorization [3]. However, more recent works like DCLM and FineWeb-edu use sharded deduplication for engineering and performance reasons. TxT360 [17] takes a slightly different approach, using global deduplication followed by upsampling by the natural distribution. We build on existing deduplication methods, and move toward upsampling based on quality metrics on a document level.
@@ 3 Repeating Filtered Datasets
In parallel to LLMs training on larger datasets, researchers are also focusing on filtering data to improve quality. Together, these factors accelerate concerns about running out of training data. In the earlier days of training LLMs, common practice was to train LLMs on a single epoch of data. This is in stark contrast to state-of-the-art vision models, where practitioners would repeat data for many epochs. Though single epoch training may have benefits such as preventing memorization [13, 3], in this work we study data repetition across datasets from a purely performance perspective. Muennighoff et al. [19] studied the repeatability of datasets in a data-constrained setting and found that datasets can be repeated for four epochs before noticing diminishing returns compared to training on fresh data. Building off this, we examine datasets of varying quality created through data filtering to understand whether these data filtering approaches affect repeatability. In this section we study datasets “as is” and treat them as a single unit rather than a collection of documents. We use original C4, DCLM’s reproduction of RefinedWeb (we will refer to this as RefinedWeb and RW v2 interchangeably), and DCLM-baseline. See Appendix I to compare datasets “as is” with datasets that are deduplicated.
@@
In Figure 2 we study a scenario similar to the findings in Muennighoff et al. [19], but instead start with Chinchilla optimal unique tokens for 1B parameter models. On the left side, we observe that the C4 validation perplexity eventually sees diminishing returns when training on C4 after repeating. In addition, we also train on higher quality datasets DCLM-baseline and RefinedWeb, and find that they have similar repeatability behavior when evaluating on C4 validation perplexity. In this setting we start with a Chinchilla optimal compute allocation and then repeat epochs by fixing the model size while increasing the token count, and compare this to datasets with more available tokens. On the right side we vary model size, but maintain the same dataset sizes that we repeat. This figure suggests that increasing model size while maintaining dataset size decreases repeatability, but there is not enough information to make a conclusion so we will explore this further later in the section. When we move to evaluating downstream accuracy across a 22 task subset from LLM Foundry (referred to as core centered metric in DCLM), we see in Figure 3 that quantifying repeatability is even less clear. It appears that when comparing using downstream accuracy, repeatability is more difficult to measure, and it is unclear if repeating meets the same diminishing returns seen when evaluating C4 perplexity. What is common between the two evaluation settings is that datasets that achieve better performance do not appear to be more repeatable than inferior datasets when examining diminishing returns.
Although better performing datasets are not more repeatable by preventing diminishing returns, in the overtraining setting (Figure 3) better datasets repeated multiple times clearly maintain better performance than an inferior dataset that is not repeated. However, it is important to note that overtraining is practical only in certain settings with priorities such as minimizing the inference cost, and there is an assumption in Figure 3 that there are at least enough tokens to train Chinchilla optimal. This assumption does not always hold when LLMs are scaled up, especially when data filtering is involved.
@@
In contrast to repeating datasets by training for longer, we next examine fixing the token budget and repeating by decreasing the number of unique tokens. This scenario forces us to examine cases where there are fewer unique tokens available than Chinchilla optimal. In Figure 4 we compare for a fixed model size and total token budget how performance varies as you reduce the number of unique tokens available, causing an increase in the number of repeats for which the model sees the dataset. When there is no longer at least the Chinchilla optimal number of tokens available, repeating the DCLM-baseline dataset with too few unique tokens can lead to worse performance than training for a single epoch on its superset of RefinedWeb filtered data. On the other hand, compute allocations with high tokens per parameter seem to benefit more from higher quality datasets, since DCLM can be repeated more times to equalize performance with unrepeated RefinedWeb. Tokens per parameter is the main factor, as increasing model size while maintaining token count decreases DCLM repeatability to equalize with RefinedWeb, and increasing token count while maintaining tokens per parameter does not consistently increase DCLM repeatability to equalize performance with unrepeated RefinedWeb (Figure1, left). This would also explain why unrepeated RefinedWeb does worse than heavily repeated DCLM in Figure3. But there are increasing diminishing returns to performance with respect to compute when increasing tokens per parameter, so there still needs to be a balance when doing data filtering.
@@
We can mitigate performance degradation caused by repeating data with regularization in the form of increasing weight decay. In Table 1, we see that the optimal weight decay value increases as we increase the number of repeats. A reasonable schedule is to scale weight decay by roughly the square root of the number of repeats. With the optimal weight decay values, we see in Figure 1 that training for ten epochs of DCLM-baseline outperforms training on RefinedWeb filtered data for a single epoch. This is notable because DCLM-baseline takes the top 10% of documents of RefinedWeb as scored by a FastText classifier. However, it is important to note that this does not mean DCLM-baseline is a strictly better dataset than RefinedWeb filtering. When data is a much more constrained resource than compute, it is likely that diminishing returns from heavily repeating DCLM-baseline will perform worse than repeating RefinedWeb filtered data a couple of times.
@@ 4 Repeating Documents To Create Better Datasets
Duplicate-aware subsample 34.1 25.1
@@
We have established that, even as compute budgets increase, models can remain robust to dataset repetition given the right training recipe. We now investigate whether high quality data repetition may be best implemented at a document level, as opposed to increased epochs on the entire dataset. To better understand this question, we dive into DCLM-baseline, which contains many duplicates (fuzzy and exact) because it uses sharded deduplication due to engineering constraints. Penedo et al. [22] also uses sharded deduplication, observing that global deduplication does not always improve performance. This is in contrast to prior standard practices of deduplication, which recommended global deduplication and training for a single epoch on the resulting dataset.
The performance gains from data filtering clearly demonstrate that not all documents are of equal quality. We can also observe this by comparing subsampling strategies. The standard strategy is uniform subsampling, which results in a dataset that is diverse with very few duplicates, while more likely keeping documents that have more duplicates. A second strategy is to run minhash-based global fuzzy deduplication before subsampling down to the desired token count. Compared to uniform subsampling, incorporating global fuzzy deduplication treats documents with different duplicate counts equally, but both strategies should have few or no duplicates in the resulting subset if subsampling significantly. Lastly, we can mimic the duplication profile of the original dataset by using duplicate-aware subsampling. For each group of duplicates, this strategy keeps or removes the entire group with probability equivalent to the duplicate count normalized by the desired dataset reduction. This strategy results in a subset with a significant number of duplicates compared to the other subsampling strategies.
@@
Figure 5: Examples of strategies for count manipulation. Note that Greedy 1 copy is similar to global deduplication, and can result in worse performance if document quality varies significantly. For example, if the dataset is split evenly between high quality documents and low quality documents, it may be better to repeat the high quality documents.
In Table 2 we observe that duplicate-aware subsampling performs the worst, suggesting that keeping a significant number of duplicates hurts performance. Uniform subsampling does significantly better than global deduplication before subsampling, which shows that documents in DCLM-baseline with high duplicate count are likely to be of higher quality. This is further supported by the fact that keeping unique copies for documents with greater than seven duplicates achieves competitive performance to uniform; however, this removes over 96% of the documents in DCLM-baseline.
@@
Table 3: 7B models trained for 138B tokens on DCLM-baseline with different document count manipulation strategies. Figure 5 visualizes these strategies. Ensemble uses the worst rank between FastText score and duplication count.
@@
Table 4: 7B models trained for 138B tokens by doing count manipulation on a duplicate-aware subsampled 280B tokens from DCLM-baseline. In this setting, the greedy yet diverse strategy sees each document approximately 3 times. Metric used is FastText score. Note that using duplication count as the metric would be ineffective here because there are not enough unique documents.
@@
The results in Table 3 are unusual in that we are taking an already highly filtered dataset and subsampling even further. Naturally, there are enough high quality tokens to train without repeats, hence the success of the greedy yet diverse strategy. However, this becomes impractical as we scale and there are no longer enough unique tokens to support this strategy. Therefore we now investigate the setting where the dataset we are trying to manipulate is closer in size to the desired token count. We duplicate-aware subsample DCLM-baseline down to 280B tokens and apply count manipulation to produce 138B tokens for training a 7B model. As we scale back up, this would allow us to use the strategy on the entire DCLM-baseline to create a dataset close to two trillion tokens. In Table 4, we see that the best strategy is applying a linear function to balance the tradeoff between exploring diverse documents and exploiting high quality documents. Contrasting the best strategy in the two settings suggests that using a good count manipulation strategy is more important in the setting where data is constrained.
@@
Combining duplicate-aware subsampling with count manipulation is a powerful tool for studying datasets. Increasing dataset size typically refers to obtaining new tokens of the same quality. DCLM does this by increasing the number of Common Crawl WARC files at the same rate as increasing the training token budget. However, as seen in Figure 6, when the number of WARC files increases, so does the number of duplicates. We find that DCLM-baseline constructed from the DCLM 7B_2x pool contains 33% fuzzy duplicates, but DCLM-baseline constructed from 20 times more WARC files contains 83% fuzzy duplicates. By uniformly subsampling to run smaller scale experiments, this experiment design uses both new data and repetitions when scaling back up. Additionally, it is impractical to deduplicate unfiltered data due to engineering constraints caused by the amount of data, as well as design decisions like what type of deduplication to use. Duplicate-aware subsampling gives us the tools to design experiments where data scaling happens through providing new data of equal quality.